1.Valles de Benasque y Eriste

TRACK OF THE ’BALL’ (3 stages) Total time: 5 hours (approx.)

a. From ’Benasque’ to Los ’Baños’ Time: 1 hour 50 min.

Depart from Benasque to the North by the road to Los Baños. After 100 m. take a track to the left ( between the river and the road). You will find a bridge over the river Ésera. After crossing it, follow the track on the left bank of the river, which is, again, the track of the Ball. Follow the tiara of the old paved track, which ends up at the main road. Keep walking and after approximately 200 m. turn left and you will soon get to a deviation.

Keep walking following the left bank of the river and in 7 minutes you will reach the old Cuera bridge. \underline{Do not cross it}, but keep walking straight and you will pass near the Aneto Camping on your way to the San Jaime bridge (this bit coincides with the GR 11). Walk under it.

After a while you will reach a trail, follow it on its left, cross the Rainero ravine, leave the descent of the Ixeira Camping on your left, and go up following the riverbed against the flow. After about 18 minutes from the Cuera bridge, the trail ascends in a bend and after a few meters it has to be left to take a path. Take it and after five minutes, you will go back to the trail and continue to the left to begin the descent between the wood and the surroundings of the Paso Nuevo dam.

Later on cross the Vallibierna river and join another trail. Here you leave the GR 11 (”Gran Recorrido” or Long Distance path) that goes to Vallibierna. Keep walking through the path on your left to the camping area of Plan de Senasta. Cross the Canal Seca (Dry Stream) and in 15 minutes you will get to Plan de Rosec, near the ruins of an old Civil Guard post where you begin the ascent again.

On your left you will see the San Farré waterfall and after about 7 min., the fountain of the same name on your right. Once passed the Cregueña and the Ésera rivers, the fishermen refuge, and the Plan de l’Acampamén, you will hear on your right the Caseta del Rey waterfall and will get to a fork. Take the track on the right and in approximately 7 min. you will reach the Plan de Baños. Cross the bridge and after a few metres take the path on your right to get to the road to Los Baños.

The altitude of the spot where Los Baños (or Bañs) is placed is 1.702 m. and they are the highest ones in Spain. It is believed that these thermal waters were already used in Roman times by the inhabitants of the valley. There are various springs, each of them at different temperature, used as part of therapeutic treatments. The main structure of the building was constructed in 1801 and consists of three floors. Los Baños were destroyed after the Spanish Civil War and later re-built. b. From ’Los Baños’ to the ’Hospital’ Time: 1 hour

Follow a forest trail that departs from the Thermal Baths and goes down to the bottom of the valley. Parallel to it there is a wide track which crosses some gravel, goes under the Turonet de Alba and continues close to the mountain slope until it reaches a point above the Pllau del Ospital. During this last bit, the track descends to the asphalt trail that goes to the Benasque Hospital. Cross this trail, keep on walking for five minutes through the prairie, by the margin of the river till you get to the Ospital de Benàs.

The site of the Hospital is at 1.758 m. It used to be a rest and shelter place for hikers and travellers. Very close to the most frequented mountain ports, it was built and protected by various religious orders, being some of the most important ones the ”Templarios” and ”Los Caballeros de San Juan en Jerusalén”.

The Benasque Hospital is known to have existed since the XIIth Century. Pilgrims, trademen and travellers were its most frequent guests. It has been destroyed several times and been used for different purposes.

c. From the ’Hospital’ to the ’Renclusa’ Time: 2 hour 20 min.

Cross the Ésera and follow the right bank towards a stony area. Keep walking following the stone wall till you get to the end of the Llanos del Hospital. Then go up the river (against the flow) following the Canal de Peña Blanca and after a few metres take the ascending path through a black pine wood to the Puyadeta de les Araneses, water pools area. (25 min. altogether).

Ten minutes later, you will get to the plains and ibones (mountain little lakes) of Plan d’Están, circumvent them on their left till you get to a higher track. Follow it and you will end up in a road behind the refuge of Plan d’Están and a little ascent that will take you to Besurta in 15 min.

When you get to the end of the road, you will find a fork with two possible routes:

- the one on the left will take you to the bottom of the valley and to the Forau d’Aigualluts,

- the one on the right ascends to the Renclusa refuge under the peak of the same name. You will get up there in approx. 45 min. This is the starting point to ascend the Aneto (3404m).